# Can LLMs Reason Structurally? An Evaluation via the Lens of Data Structures

### v2

@@ A.4.5 Performance of DeepSeek-R1
Associative Hashmap Compound 0.44 (0.05) 0.01 (0.02) 0.03 (0.03)
Trie* Compound 0.54 (0.24) 0.32 (0.04) 0.12 (0.06)
Suffix Tree* Construct 0.93 (0.07) 0.50 (0.07) 0.05 (0.05)
@@
A.4.6 Performance of GPT-4.1
Table 18: Mean (± std) accuracy of GPT-4.1 on all DSR-Bench tasks over three runs. Data structures marked with * are included in the DSR-Bench-challenge suite.
@@
Associative Hashmap Compound 0.19 (0.07) 0.00 (0.00) 0.00 (0.00)
Trie* Compound 0.39 (0.07) 0.13 (0.03) 0.01 (0.02)
Suffix Tree* Construct 0.00 (0.00) 0.00 (0.00) 0.00 (0.00)
@@
A.4.7 Performance of Gemini-2.0-Flash
Table 19: Mean (± std) accuracy of Gemini-2.0-Flash on all DSR-Bench tasks over three runs. Data structures marked with * are included in the DSR-Bench-challenge suite.
@@
Associative Hashmap Compound 0.28 (0.05) 0.01 (0.02) 0.00 (0.00)
Trie* Compound 0.31 (0.02) 0.18 (0.02) 0.03 (0.00)
Suffix Tree* Construct 0.00 (0.00) 0.02 (0.02) 0.00 (0.00)
@@
A.4.8 Performance of Claude-3.5-Sonnet
Table 20: Mean (± std) accuracy of Claude-3.5-Sonnet on all DSR-Bench tasks over three runs. Data structures marked with * are included in the DSR-Bench-challenge suite.
@@
A.4.9 Performance of DeepSeek-V3
Table 21: Mean (± std) accuracy of DeepSeek-V3 on all DSR-Bench tasks over three runs. Data structures marked with * are included in the DSR-Bench-challenge suite.
@@ A.4.9 Performance of DeepSeek-V3
Associative Hashmap Compound 0.04 (0.02) 0.00 (0.00) 0.00 (0.00)
Trie* Compound 0.00 (0.00) 0.00 (0.00) 0.00 (0.00)
Suffix Tree* Construct 0.06 (0.02) 0.00 (0.00) 0.00 (0.00)
@@
A.4.10 Performance of Llama-3.3
Table 22: Mean (± std) accuracy of Llama-3.3 on all DSR-Bench tasks over three runs. Data structures marked with * are included in the DSR-Bench-challenge suite.
@@
Associative Hashmap Compound 0.00 (0.00) 0.00 (0.00) 0.00 (0.00)
Trie* Compound 0.01 (0.02) 0.00 (0.00) 0.00 (0.00)
Suffix Tree* Construct 0.00 (0.00) 0.00 (0.00) 0.00 (0.00)
@@
A.5 Accuracy by prompting methods across instruction-tuned models
@@
This section presents additional accuracy tables for tasks in DSR-Bench, evaluating each instruction-tuned model across five prompting methods: Stepwise, 0-CoT, CoT, 3-shot, and None. The results are shown in Table˜23 (GPT-4.1), Table˜24 (Gemini-2.0-Flash), Table˜25 (Claude-3.5-Sonnet), Table˜26 (DeepSeek-V3), and Table˜27 (Llama-3.3).
@@
A.6 Additional analysis on CoT prompting
@@
A.7 The spatial suite supplementary materials
Figure˜3 presents example illustrations of the non-uniform input distributions: circles, moons, and blobs, which were adopted from scikit-learn. These synthetic patterns are used to evaluate whether models can adapt to irregular and non-uniform spatial distributions, an essential aspect of real-world data, as discussed in Section˜4.2.
@@ A.7 The spatial suite supplementary materials
A.8 The natural suite supplementary materials
@@
Figure˜4 illustrates our generation process for natural language prompts in DSR-Bench-natural. For each data structure task, we begin by manually writing an initial scenario narrative (see On a sunny afternoon… arrived in the queue” in Section˜A.8). This narrative is then paraphrased into five variants using GPT-4o. For each supported operation type (e.g., enqueue, dequeue), we generate 30 paraphrased operation templates in a similar manner (e.g., With money in hand, Yuki Lopez stepped into the queue.” corresponds to “enqueue Yuki Lopez” in Section˜A.8). In total, each task is backed by a pool of five scenario descriptions and 30 templates per operation type.
@@
A.9 The code suite supplementary materials
Our benchmark is designed to evaluate general reasoning in LLMs without relying on external tools such as code interpreters or formal solvers. The focus is on the underlying reasoning ability that drives broad problem-solving, beyond coding or domains where code is directly applicable. This follows recent evaluation practices, such as Gemini-Deep-Think and OpenAI’s IMO assessments (Luong & Lockhart, 2025; Wei, 2025), which deliberately disallowed code or tool use (e.g., Lean) to test end-to-end reasoning. By using data structures as controlled, interpretable settings for structural reasoning—a core component of general reasoning—DSR-Bench differs from code-synthesis benchmarks that emphasise syntax or API knowledge and are often vulnerable to data leakage.
To additionally probe how code generation may support structural reasoning, we conducted an ablation across six models, including reasoning models with competitive performance on code-synthesis benchmarks. We cover seven data structures of varying difficulty (Array, Queue, Hashmap, Heap, DAWG, Geom Graph, Graph-Natural) and average results over three runs. Three code-generation modes were tested, with prompts shown below:
@@ A.9 The code suite supplementary materials
In CodeOnly, models perform well on Geom Graph (k-dimensional graphs embedded in geometric space, which is a standard data structure widely used in computer graphics), whose code implementation is more available online. However, they struggle with the less familiar DAWG” (directed acyclic word graph), where we define it with customized constraints to enforce a unique output. This suggests that performance may reflect memorization rather than true reasoning.
@@
A.10 Levenshtein distance: an auxiliary metric for DSR-Bench
@@
However, this granularity can also blur important semantic distinctions. A syntactically well-formed but semantically incorrect output may still receive a high score, especially when the expected answer is long or formatted. When averaged over 30 test cases, models with large gaps in binary accuracy can appear deceptively similar under Levenshtein. For example, the Skip List and DSU compound tasks yield the same Levenshtein score (0.75), despite the former achieving more than triple the binary accuracy (Table˜32).
@@
A.11 Failure rates of JSON parsing via Structured Output
@@
A.12 Ablation on paraphrased prompt templates
In the natural suite, we use 5 paraphrased templates for problem statements and 30 for operations (e.g., “A space tunnel connects planet1 and planet2”) to smooth out ambiguity in natural language descriptions. However, in the main and challenge suites, we use formalized prompt templates to enable large-scale evaluation. We conducted an ablation to test how prompt template variations affect performance. Prompting GPT-4o with “Paraphrase the following description,” we generated five paraphrases and found that overall performance trends remained consistent.
@@ A.12 Ablation on paraphrased prompt templates
A.13 Preliminary study on strategy adaptation with DSR-Bench
@@
Beyond implicit evaluation, DSR-Bench’s modular design—data generation, prompt templates, automated evaluation, and schema-based verification—readily supports dynamic strategy switching. A preliminary study is shown in Table˜36 and Table˜35. We mask structure names or add constraints to DSR-Bench’s prompts, and test whether models can, (i) select BFS vs DFS for graph tasks, or (ii) choose between K-D tree and array for nearest-neighbor queries under varying complexities.

### v3

@@ D.5 Performance of DeepSeek-R1
Associative Hashmap* Compound 0.44 (0.05) 0.01 (0.02) 0.03 (0.03)
Trie Compound 0.54 (0.24) 0.32 (0.04) 0.12 (0.06)
Suffix Tree Construct 0.93 (0.07) 0.50 (0.07) 0.05 (0.05)
@@
D.6 Performance of GPT-4.1
Table 12: Mean (± std) accuracy of GPT-4.1 on all DSR-Bench tasks over three runs. Data structures marked with * are included in the DSR-Bench-challenge subset.
@@
Associative Hashmap* Compound 0.19 (0.07) 0.00 (0.00) 0.00 (0.00)
Trie Compound 0.39 (0.07) 0.13 (0.03) 0.01 (0.02)
Suffix Tree Construct 0.00 (0.00) 0.00 (0.00) 0.00 (0.00)
@@
D.7 Performance of Gemini-2.0-Flash
Table 13: Mean (± std) accuracy of Gemini-2.0-Flash on all DSR-Bench tasks over three runs. Data structures marked with * are included in the DSR-Bench-challenge subset.
@@
Associative Hashmap* Compound 0.28 (0.05) 0.01 (0.02) 0.00 (0.00)
Trie Compound 0.31 (0.02) 0.18 (0.02) 0.03 (0.00)
Suffix Tree Construct 0.00 (0.00) 0.02 (0.02) 0.00 (0.00)
@@
D.8 Performance of Claude-3.5-Sonnet
Table 14: Mean (± std) accuracy of Claude-3.5-Sonnet on all DSR-Bench tasks over three runs. Data structures marked with * are included in the DSR-Bench-challenge subset.
@@
D.9 Performance of DeepSeek-V3
Table 15: Mean (± std) accuracy of DeepSeek-V3 on all DSR-Bench tasks over three runs. Data structures marked with * are included in the DSR-Bench-challenge subset.
@@ D.9 Performance of DeepSeek-V3
Associative Hashmap* Compound 0.04 (0.02) 0.00 (0.00) 0.00 (0.00)
Trie Compound 0.00 (0.00) 0.00 (0.00) 0.00 (0.00)
Suffix Tree Construct 0.06 (0.02) 0.00 (0.00) 0.00 (0.00)
@@
D.10 Performance of Llama-3.3
Table 16: Mean (± std) accuracy of Llama-3.3 on all DSR-Bench tasks over three runs. Data structures marked with * are included in the DSR-Bench-challenge subset.
@@
Associative Hashmap* Compound 0.00 (0.00) 0.00 (0.00) 0.00 (0.00)
Trie Compound 0.01 (0.02) 0.00 (0.00) 0.00 (0.00)
Suffix Tree Construct 0.00 (0.00) 0.00 (0.00) 0.00 (0.00)
@@
D.11 Performance of Qwen3-8b-Instruct
Table 17: Mean (± std) accuracy of Qwen3-8b-Instruct on all DSR-Bench tasks over three runs. Data structures marked with * are included in the DSR-Bench-challenge subset.
Category Data Structure Operation Short Medium Long
Linear Array Access 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
Delete 0.99 (0.02) 0.99 (0.02) 0.96 (0.04)
Insert 1.00 (0.00) 1.00 (0.00) 0.97 (0.03)
Reverse 1.00 (0.00) 1.00 (0.00) 0.92 (0.02)
Search 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
Temporal Stack Compound 1.00 (0.00) 0.99 (0.02) 0.93 (0.06)
Queue Compound 1.00 (0.00) 0.98 (0.04) 0.98 (0.02)
LRU Cache Compound 0.93 (0.00) 0.77 (0.09) 0.49 (0.08)
Priority Queue* Compound 0.71 (0.07) 0.28 (0.08) 0.07 (0.03)
Associative Hashmap* Compound 0.59 (0.08) 0.68 (0.10) 0.20 (0.00)
Trie Compound 0.02 (0.02) 0.00 (0.00) 0.01 (0.02)
Suffix Tree Construct 0.00 (0.00) 0.00 (0.00) 0.00 (0.00)
Skip List* Compound 0.08 (0.05) 0.00 (0.00) 0.00 (0.00)
Hierarchical BST Insert 0.78 (0.07) 0.30 (0.13) 0.01 (0.02)
Remove 0.63 (0.06) 0.26 (0.04) 0.01 (0.02)
In-Order Traversal 0.84 (0.02) 0.57 (0.07) 0.23 (0.03)
Pre-Order Traversal 0.87 (0.06) 0.80 (0.09) 0.84 (0.02)
Post-Order Traversal 0.86 (0.12) 0.76 (0.07) 0.37 (0.07)
Depth 1.00 (0.00) 0.97 (0.03) 0.89 (0.04)
Compound 0.87 (0.03) 0.24 (0.08) 0.01 (0.02)
Heap* Compound 0.34 (0.07) 0.07 (0.03) 0.02 (0.02)
Heapify 0.79 (0.05) 0.24 (0.04) 0.01 (0.02)
RB Tree* Construct 0.00 (0.00) 0.00 (0.00) 0.00 (0.00)
Compound 0.22 (0.02) 0.00 (0.00) 0.00 (0.00)
B+ Tree* Compound 0.23 (0.15) 0.08 (0.05) 0.03 (0.03)
K-D Tree* Compound 0.01 (0.02) 0.00 (0.00) 0.00 (0.00)
K-D Heap* Compound 0.12 (0.05) 0.03 (0.00) 0.00 (0.00)
Network Graph Depth-First Traversal 0.63 (0.03) 0.04 (0.02) 0.00 (0.00)
Breadth-First Traversal 0.46 (0.13) 0.01 (0.02) 0.00 (0.00)
DSU Compound 0.29 (0.04) 0.01 (0.02) 0.00 (0.00)
Geom Graph* Construct 0.00 (0.00) 0.00 (0.00) 0.00 (0.00)
Hybrid Bloom Filter* Compound 0.36 (0.10) 0.00 (0.00) 0.00 (0.00)
DAWG* Compound 0.01 (0.02) 0.00 (0.00) 0.00 (0.00)
D.12 Performance of Phi-4-reasoning-14B
Table 18: Mean (± std) accuracy of Phi-4-reasoning-14B on all DSR-Bench tasks over three runs. Data structures marked with * are included in the DSR-Bench-challenge subset.
Category Data Structure Operation Short Medium Long
Linear Array Access 0.97 (0.03) 0.77 (0.05) 0.73 (0.09)
Delete 0.85 (0.02) 0.65 (0.02) 0.53 (0.07)
Insert 0.85 (0.02) 0.54 (0.02) 0.24 (0.08)
Reverse 0.68 (0.07) 0.08 (0.07) 0.03 (0.03)
Search 0.97 (0.00) 0.83 (0.00) 0.70 (0.03)
Temporal Stack Compound 0.15 (0.02) 0.00 (0.00) 0.00 (0.00)
Queue Compound 0.33 (0.05) 0.02 (0.02) 0.00 (0.00)
LRU Cache Compound 0.43 (0.05) 0.02 (0.02) 0.03 (0.00)
Priority Queue* Compound 0.02 (0.02) 0.00 (0.00) 0.00 (0.00)
Associative Hashmap* Compound 0.00 (0.00) 0.00 (0.00) 0.00 (0.00)
Trie Compound 0.00 (0.00) 0.00 (0.00) 0.00 (0.00)
Suffix Tree Construct 0.00 (0.00) 0.00 (0.00) 0.00 (0.00)
Skip List* Compound 0.02 (0.02) 0.00 (0.00) 0.00 (0.00)
Hierarchical BST Insert 0.31 (0.02) 0.01 (0.02) 0.00 (0.00)
Remove 0.14 (0.05) 0.02 (0.02) 0.00 (0.00)
In-Order Traversal 0.50 (0.05) 0.37 (0.00) 0.20 (0.07)
Pre-Order Traversal 0.68 (0.07) 0.62 (0.07) 0.41 (0.07)
Post-Order Traversal 0.13 (0.14) 0.02 (0.02) 0.00 (0.00)
Depth 0.28 (0.02) 0.17 (0.09) 0.23 (0.00)
Compound 0.10 (0.05) 0.00 (0.00) 0.00 (0.00)
Heap* Compound 0.08 (0.02) 0.00 (0.00) 0.00 (0.00)
Heapify 0.10 (0.05) 0.00 (0.00) 0.00 (0.00)
RB Tree* Construct 0.00 (0.00) 0.00 (0.00) 0.00 (0.00)
Compound 0.07 (0.00) 0.00 (0.00) 0.00 (0.00)
B+ Tree* Compound 0.07 (0.00) 0.00 (0.00) 0.00 (0.00)
K-D Tree* Compound 0.00 (0.00) 0.00 (0.00) 0.00 (0.00)
K-D Heap* Compound 0.03 (0.00) 0.00 (0.00) 0.00 (0.00)
Network Graph Depth-First Traversal 0.07 (0.00) 0.01 (0.02) 0.00 (0.00)
Breadth-First Traversal 0.07 (0.00) 0.00 (0.00) 0.00 (0.00)
DSU Compound 0.00 (0.00) 0.00 (0.00) 0.00 (0.00)
Geom Graph* Construct 0.00 (0.00) 0.00 (0.00) 0.00 (0.00)
Hybrid Bloom Filter* Compound 0.02 (0.02) 0.00 (0.00) 0.00 (0.00)
DAWG* Compound 0.00 (0.00) 0.00 (0.00) 0.00 (0.00)
D.13 Performance of Mixtral-8x7B
Table 19: Mean (± std) accuracy of Mixtral-8x7B on all DSR-Bench tasks over three runs. Data structures marked with * are included in the DSR-Bench-challenge subset.
Category Data Structure Operation Short Medium Long
Linear Array Access 0.87 (0.00) 0.55 (0.02) 0.26 (0.04)
Delete 0.88 (0.02) 0.54 (0.05) 0.30 (0.00)
Insert 0.68 (0.02) 0.63 (0.05) 0.22 (0.04)
Reverse 0.60 (0.03) 0.03 (0.00) 0.02 (0.02)
Search 0.87 (0.00) 0.48 (0.02) 0.38 (0.02)
Temporal Stack Compound 0.08 (0.02) 0.00 (0.00) 0.00 (0.00)
Queue Compound 0.26 (0.02) 0.01 (0.02) 0.00 (0.00)
LRU Cache Compound 0.04 (0.05) 0.02 (0.02) 0.00 (0.00)
Priority Queue* Compound 0.06 (0.02) 0.00 (0.00) 0.00 (0.00)
Associative Hashmap* Compound 0.00 (0.00) 0.00 (0.00) 0.00 (0.00)
Trie Compound 0.00 (0.00) 0.02 (0.02) 0.00 (0.00)
Suffix Tree Construct 0.00 (0.00) 0.00 (0.00) 0.00 (0.00)
Skip List* Compound 0.00 (0.00) 0.00 (0.00) 0.00 (0.00)
Hierarchical BST Insert 0.00 (0.00) 0.00 (0.00) 0.00 (0.00)
Remove 0.00 (0.00) 0.00 (0.00) 0.00 (0.00)
In-Order Traversal 0.45 (0.02) 0.30 (0.05) 0.11 (0.02)
Pre-Order Traversal 0.33 (0.00) 0.10 (0.03) 0.07 (0.00)
Post-Order Traversal 0.04 (0.02) 0.00 (0.00) 0.00 (0.00)
Depth 0.28 (0.02) 0.05 (0.02) 0.00 (0.00)
Compound 0.11 (0.02) 0.00 (0.00) 0.00 (0.00)
Heap* Compound 0.00 (0.00) 0.00 (0.00) 0.00 (0.00)
Heapify 0.04 (0.02) 0.00 (0.00) 0.00 (0.00)
RB Tree* Construct 0.00 (0.00) 0.00 (0.00) 0.00 (0.00)
Compound 0.00 (0.00) 0.00 (0.00) 0.00 (0.00)
B+ Tree* Compound 0.03 (0.00) 0.00 (0.00) 0.00 (0.00)
K-D Tree* Compound 0.00 (0.00) 0.00 (0.00) 0.00 (0.00)
K-D Heap* Compound 0.03 (0.00) 0.00 (0.00) 0.00 (0.00)
Network Graph Depth-First Traversal 0.00 (0.00) 0.03 (0.00) 0.00 (0.00)
Breadth-First Traversal 0.03 (0.00) 0.00 (0.00) 0.00 (0.00)
DSU Compound 0.00 (0.00) 0.00 (0.00) 0.00 (0.00)
Geom Graph* Construct 0.00 (0.00) 0.00 (0.00) 0.00 (0.00)
Hybrid Bloom Filter* Compound 0.00 (0.00) 0.00 (0.00) 0.00 (0.00)
DAWG* Compound 0.00 (0.00) 0.00 (0.00) 0.00 (0.00)
@@ D.13 Performance of Mixtral-8x7B
Appendix E Accuracy by prompting methods across instruction-tuned models
@@
This section presents additional accuracy tables for tasks in DSR-Bench, evaluating each instruction-tuned model across five prompting methods: Stepwise, 0-CoT, CoT, 3-shot, and None. The results are shown in Table˜20 (GPT-4.1), Table˜21 (Gemini-2.0-Flash), Table˜22 (Claude-3.5-Sonnet), Table˜23 (DeepSeek-V3), and Table˜24 (Llama-3.3).
@@
E.1 Additional analysis on CoT prompting
@@
Appendix F The spatial probe supplementary materials
Figure˜8 presents example illustrations of the non-uniform input distributions: circles, moons, and blobs, which were adopted from scikit-learn. These synthetic patterns are used to evaluate whether models can adapt to irregular and non-uniform spatial distributions, an essential aspect of real-world data, as discussed in Section˜4.2.
@@ Appendix F The spatial probe supplementary materials
Appendix G The realistic probe supplementary materials
@@
Figure˜9 illustrates our generation process for natural language prompts in DSR-Bench-realistic. For each data structure task, we begin by manually writing an initial scenario narrative (see On a sunny afternoon… arrived in the queue” in Appendix˜G). This narrative is then paraphrased into five variants using GPT-4o. For each supported operation type (e.g., enqueue, dequeue), we generate 30 paraphrased operation templates in a similar manner (e.g., With money in hand, Yuki Lopez stepped into the queue.” corresponds to “enqueue Yuki Lopez” in Appendix˜G). In total, each task is backed by a pool of five scenario descriptions and 30 templates per operation type.
@@
Appendix H The code probe supplementary materials
Our benchmark is designed to evaluate general (inherent) reasoning in LLMs without relying on external tools such as code interpreters or formal solvers. The focus is on the underlying reasoning ability that drives broad problem-solving, beyond coding or domains where code is directly applicable. This follows recent evaluation practices, such as Gemini-Deep-Think and OpenAI’s IMO assessments (Luong and Lockhart, 2025; Wei, 2025), which deliberately disallowed code or tool use (e.g., Lean) to test end-to-end reasoning. By using data structures as controlled, interpretable settings for structural reasoning, DSR-Bench differs from code-synthesis benchmarks that emphasize syntax or API knowledge and are often vulnerable to training data contamination.
To further investigate how code generation may support structural reasoning, we conducted an ablation across six models, including reasoning models with competitive performance on code-synthesis benchmarks. We cover seven data structures of varying difficulty (Array, Queue, Hashmap, Heap, DAWG, Geom Graph, Graph-Natural) and average results over three runs. Three code-generation modes were tested, with prompts shown below:
@@ Appendix H The code probe supplementary materials
In CodeOnly, models perform well on Geom Graph (k-dimensional graphs embedded in geometric space, which is a standard data structure widely used in computer graphics), whose code implementation is more available online. However, they struggle with the less familiar DAWG (directed acyclic word graph), where we define it with customized constraints to enforce a unique output. This suggests that performance may reflect memorization rather than true reasoning.
@@
Appendix I Levenshtein distance: an auxiliary metric for DSR-Bench
@@
However, this granularity can also blur important semantic distinctions. A syntactically well-formed but semantically incorrect output may still receive a high score, especially when the expected answer is long or formatted. When averaged over 30 test cases, models with large gaps in binary accuracy can appear deceptively similar under Levenshtein. For example, the Skip List and DSU compound tasks yield the same Levenshtein score (0.75), despite the former achieving more than triple the binary accuracy (Table˜29).
@@
Appendix J Failure rates of JSON parsing via Structured Output
@@
Appendix K Ablation on paraphrased prompt templates
In the realistic probe, we use 5 paraphrased templates for problem statements and 30 for operations (e.g., “A space tunnel connects planet1 and planet2”) to smooth out ambiguity in natural language descriptions. However, in the main and challenge suites, we use formalized prompt templates to enable large-scale evaluation. We conducted an ablation to test how prompt template variations affect performance. Prompting GPT-4o with “Paraphrase the following description,” we generated five paraphrases and found that overall performance trends remained consistent.
@@
Appendix L Preliminary study on strategy adaptation with DSR-Bench
@@
Beyond implicit evaluation, DSR-Bench’s modular design (data generation, prompt templates, automated evaluation, and schema-based verification) readily supports dynamic strategy switching. A preliminary study is shown in Table˜33 and Table˜32. We mask structure names or add constraints to DSR-Bench’s prompts, and test whether models can, (i) select BFS vs DFS for graph tasks, or (ii) choose between K-D tree and array for nearest-neighbor queries under varying complexities.